# MEMOIR: Lifelong Model Editing with Minimal Overwrite and Informed Retention for LLMs

### v3

@@ 4.1 Experimental setup
We compare MEMOIR against several established baselines. We first consider non-parametric methods, namely GRACE grace , which stores knowledge in an external codebook. We also evaluate parametric methods, including DEFER grace , a method inspired by SERAC serac for inference-time routing; methods based on causal tracing, including ROME rome , MEMIT memit , and AlphaEdit fang2025alphaedit ; and memory-based approaches such as WISE wang2024wise . Finally, we include direct fine-tuning (FT) as an additional baseline. Detailed descriptions and implementation details for all methods are provided in Appendix A.
@@
We perform our experiments on four different widely used autoregressive LLMs: LLaMA-3-8B-Instruct llama-3 , Mistral-7Bmistral , LLaMA-2-7Bllama-2 (results provided in Appendix B.1) and GPT-J-6B gpt-j . For brevity, we refer to them as LLaMA-3, Mistral, LLaMA-2, and GPT-J throughout this section.
@@
ROME rome 0.99 0.96 0.96 0.97 0.61 0.59 0.41 0.54 0.08 0.08 0.02 0.06 0.03 0.03 0.02 0.03
MEMIT memit 0.99 0.97 0.98 0.98 0.68 0.67 0.73 0.69 0.03 0.03 0.01 0.02 0.00 0.00 0.00 0.00
DEFER grace 0.82 0.86 0.68 0.79 0.66 0.67 0.45 0.59 0.33 0.32 1.00 0.55 0.31 0.31 1.00 0.54
GRACE grace 1.00 0.46 1.00 0.82 1.00 0.42 1.00 0.81 1.00 0.39 1.00 0.80 1.00 0.37 1.00 0.79
WISE wang2024wise 0.92 0.84 1.00 0.92 0.78 0.74 0.98 0.83 0.62 0.60 1.00 0.74 0.66 0.64 1.00 0.77
AlphaEdit fang2025alphaedit 0.98 0.89 1.00 0.96 0.93 0.85 0.98 0.92 0.91 0.79 0.94 0.88 0.84 0.77 0.56 0.72
@@
ROME rome 0.79 0.77 0.98 0.85 0.58 0.57 0.75 0.63 0.05 0.05 0.02 0.04 0.04 0.04 0.02 0.03
MEMIT memit 0.81 0.79 0.99 0.86 0.46 0.45 0.61 0.51 0.00 0.00 0.01 0.00 0.04 0.04 0.02 0.03
DEFER grace 0.59 0.68 0.79 0.69 0.44 0.43 1.00 0.62 0.40 0.40 1.00 0.60 0.40 0.39 1.00 0.60
GRACE grace 1.00 0.36 1.00 0.79 1.00 0.15 1.00 0.72 1.00 0.15 1.00 0.72 1.00 0.02 1.00 0.67
WISE wang2024wise 0.98 0.97 1.00 0.98 0.92 0.89 1.00 0.94 0.87 0.80 1.00 0.89 0.70 0.67 1.00 0.79
AlphaEdit fang2025alphaedit 0.83 0.77 1.00 0.87 0.87 0.75 0.99 0.87 0.86 0.74 0.95 0.85 0.85 0.72 0.68 0.75
@@
ROME rome 1.52 0.97 2.02e1 0.61 1.26e5 0.02 4.07e4 0.02     2.04 0.99 3.45 0.92 1.04e2 0.03 2.41e2 0.01
MEMIT memit 1.30 0.99 4.05e2 0.93 4.56e10 0.01 2.95e6 0.00     1.64 1.00 1.59e1 0.89 9.72e1 0.04 1.32e2 0.02
DEFER grace 5.69e2 0.83 2.36e2 0.82 7.02e1 1.00 6.85e1 1.00     1.46e1 0.94 3.22e1 0.99 3.22e1 0.99 3.22e1 1.00
GRACE grace 1.05 1.00 7.10e1 1.00 7.12e1 1.00 7.73e1 1.00     1.39 1.00 5.97 1.00 9.53 1.00 9.57 1.00
WISE wang2024wise 4.93e1 0.98 1.46 0.95 2.10 0.99 3.20 0.99     1.40 1.00 2.56 0.94 1.31 0.99 5.21 0.93
AlphaEdit fang2025alphaedit 1.58 1.00 3.12 0.98 5.97 0.93 8.49e3 0.05     1.75 1.00 1.76 1.00 2.87 0.98 1.70e2 0.88
@@
We follow prior work and evaluate the effectiveness of different editing methods based on three metrics (wang2024wise, ): reliability, generalization and localization, which are respectively computed by model’s performance on edit, rephrased edit and irrelevant samples.
@@ 4.2 Main Experimental Results
Table 1 presents the results on the ZsRE zsre dataset for the question answering task, evaluated across a range of total edits from $1$ to $1000$. ZsRE is a context-free Q&A benchmark designed to assess a model’s ability to store and retrieve specific factual knowledge. We observe that MEMOIR consistently achieves the most balanced performance across all evaluation settings. On LLaMA-3, MEMOIR maintains a reliability of 0.94 and a generalization of 0.85, with its locality reaching 1.0 across different edit counts. In contrast, other parametric methods exhibit substantial degradation as the number of edits increases. For example, WISE drops from an average metric of 0.92 to 0.77 as the number of edits grows from $1$ to 1,0001{,}000. A similar trend is observed with AlphaEdit, which declines from 0.96 to 0.72. Although GRACE, a non-parametric baseline, maintains high reliability, it suffers from poor generalization, leading to a significantly lower average metric overall compared to MEMOIR. In summary, MEMOIR offers the best trade-off between reliability, generalization, and locality, achieving an average metric of 0.93 on LLaMA-3 with 1,0001{,}000 total edits—outperforming all prior methods by a substantial margin of 0.14. Similar results are observed with Mistral, where MEMOIR again delivers the highest average metric, demonstrating its robustness and effectiveness across different LLMs.
@@
LLMs are prone to hallucinations—generating fluent but factually inaccurate or misleading content—which poses a significant challenge for their deployment in domains that demand high factual precision. We validate here the effectiveness of MEMOIR on correcting hallucinations on the SelfCheckGPT selfcheckgpt dataset, which consists of Wikipedia-style biographies generated by GPT-3 gpt-3 . The dataset is annotated for factual hallucinations, with ground-truth replacements verified against Wikipedia. We note that we report only locality and reliability using perplexity (PPL) metric, following (wang2024wise, ), as no suitable metric for generalization exists in this setting. As shown in Table 2, MEMOIR again achieves the best balanced performance, with its advantages becoming even more significant with a larger number of edits. Under the most challenging setting of $600$ edits, MEMOIR maintains a saturated locality while delivering a perplexity metric, 61% and by 77% smaller than WISE, the second-best performing approach, respectively on LLaMA-3 and on Mistral. These results further confirm MEMOIR’s robustness in sustaining model quality during long editing sequences.
@@
MEMIT memit 0.88 0.28 0.98 0.71 0.41 0.18 0.14 0.24
GRACE grace 0.97 0.28 1.00 0.75 0.97 0.28 1.00 0.75
WISE wang2024wise 0.99 0.36 0.98 0.78 0.96 0.37 1.00 0.78
AlphaEdit fang2025alphaedit 0.89 0.27 0.99 0.72 0.85 0.27 0.96 0.69
@@
Effective model editing should generalize from structured prompts to distributionally shifted inputs, where the shift reflects increased complexity rather than domain change (domain_shift, ; canonical_examples, ). We evaluate the performance of MEMOIR on OOD settings using the Temporal dataset redpajama , which consists of a prefix paired with a Wikipedia description and a model-generated description. As shown in Table 3, MEMOIR outperforms the state-of-the-art consistently at both $T=10$ and $T=75$, indicating strong adaptation to novel knowledge and effective generalization.
@@
In appendix B.5 we assess the efficacy of MEMOIR in multi-hop reasoning, a more challenging scenario. We evaluate on the RippleEdits benchmark (ripple, ), where MEMOIR demonstrates better multi-hop reasoning generalization than all baseline methods.
@@ 4.3 Scaling with longer sequences of edits
MEMIT (memit, ) 0.00 0.00 0.00 0.00
GRACE (grace, ) 1.00 0.31 1.00 0.77 0.99 0.28 1.00 0.76
WISE (wang2024wise, ) 0.59 0.57 1.00 0.72 0.43 0.41 1.00 0.61
AlphaEdit (fang2025alphaedit, ) 0.02 0.02 0.00 0.01 0.02 0.02 0.00 0.01
@@ 6 Related Works
Sparsity and non-overlapping activations have been widely explored in continual learning to mitigate catastrophic forgetting by localizing updates and reducing interference. Early works french1991using ; french1999catastrophic introduced this idea, later extended by methods like PackNet mallya2018packnet and Supermasks-in-Superposition wortsman2020supermasks , which allocate disjoint parameter subsets per task but require task identities at inference and assume clear task boundaries. Gradient-based methods such as GPM saha2021gradient and SPARCL wang2022sparcl improve parameter efficiency via orthogonal updates or activation sparsity. Even standard Dropout srivastava_dropout_2014 implicitly encourages non-overlapping activations mirzadeh2020understanding , while heterogeneous dropout across tasks aids activation sparsity abbasi2022sparsity . However, these approaches remain confined to continual learning, often relying on irrelevant data or explicit regularization. In contrast, we tackle lifelong model editing without task supervision or boundaries, inducing non-overlapping updates by directly sparsifying input activations to a residual memory, enabling scalable, localized knowledge editing.
@@
Prior work in model editing falls into two categories: parametric and non-parametric methods. Parametric methods typically modify model weights through meta-learning, locating-then-editing strategies, or auxiliary memory modules. Meta-learning approaches like KE (ike, ) and MEND (mend, ) use hypernetworks to incorporate updates while minimizing disruption to previous knowledge. Locating-then-editing methods such as ROME (rome, ) and MEMIT (memit, ) identify and edit specific modules where facts are stored, while AlphaEdit (fang2025alphaedit, ) improves them by projecting edits onto the null space of preserved knowledge to avoid corrupting previous information. Auxiliary module methods, including SERAC (serac, ) and T-Patcher (t-patcher, ), extend the model with lightweight components for storing and routing updates. WISE (wang2024wise, ) further advances this line by introducing routing mechanisms and memory sharding to localize edits and mitigate cross-edit interference. Despite their effectiveness, parametric methods often suffer from forgetting over long edit sequences. Non-parametric methods store knowledge externally to preserve original weights. GRACE (grace, ) retrieves discrete codebook entries to overwrite intermediate activations during inference to enable precise and local edits, while LOKA (zhang2025resolving, ) tackles simultaneous editing and unlearning by storing updates in task-specific or shared memory. However, these methods rely on exact input matches, limiting generalization to paraphrased queries.
@@ 7 Conclusion and Limitations
We introduced MEMOIR, a scalable framework for lifelong model editing that balances reliability, generalization, and locality. By sparsifying input activations through sample-dependent masks and confining updates to a dedicated parameter space, MEMOIR enables precise knowledge injection without disrupting the pre-trained model’s capabilities. At inference, it retrieves relevant updates by comparing sparse activation patterns, allowing edits to generalize to rephrased queries while preserving the model’s behavior on unrelated prompts. Extensive experiments show that MEMOIR outperforms state-of-the-art methods across question answering, hallucination correction, and out-of-distribution generalization tasks, scaling to thousands of edits with minimal forgetting. Overall, MEMOIR offers a practical tool for improving the factual accuracy and safety of deployed language models [^2] Code will be released upon acceptance..

### v4

@@ 4.1 Experimental setup
We compare MEMOIR against several established baselines. We first consider non-parametric methods, namely GRACE Hartvigsen et al. (2023), which stores knowledge in an external codebook. We also evaluate parametric methods, including DEFER Hartvigsen et al. (2023), a method inspired by SERAC Mitchell et al. (2022b) for inference-time routing; methods based on causal tracing, including ROME Meng et al. (2022), MEMIT Meng et al. (2023), and AlphaEdit Fang et al. (2025); and memory-based approaches such as WISE Wang et al. (2024). Finally, we include direct fine-tuning (FT) as an additional baseline. Detailed descriptions and implementation details for all methods are provided in Appendix A.
@@
We perform our experiments on four different widely used autoregressive LLMs: LLaMA-3-8B-Instruct Aaron et al. (2024), Mistral-7BJiang et al. (2023), LLaMA-2-7BTouvron et al. (2023) (results provided in Appendix B.1) and GPT-J-6B Wang and Komatsuzaki (2021). For brevity, we refer to them as LLaMA-3, Mistral, LLaMA-2, and GPT-J throughout this section.
@@
ROME Meng et al. (2022) 0.99 0.96 0.96 0.97 0.61 0.59 0.41 0.54 0.08 0.08 0.02 0.06 0.03 0.03 0.02 0.03
MEMIT Meng et al. (2023) 0.99 0.97 0.98 0.98 0.68 0.67 0.73 0.69 0.03 0.03 0.01 0.02 0.00 0.00 0.00 0.00
DEFER Hartvigsen et al. (2023) 0.82 0.86 0.68 0.79 0.66 0.67 0.45 0.59 0.33 0.32 1.00 0.55 0.31 0.31 1.00 0.54
GRACE Hartvigsen et al. (2023) 1.00 0.46 1.00 0.82 1.00 0.42 1.00 0.81 1.00 0.39 1.00 0.80 1.00 0.37 1.00 0.79
WISE Wang et al. (2024) 0.92 0.84 1.00 0.92 0.78 0.74 0.98 0.83 0.62 0.60 1.00 0.74 0.66 0.64 1.00 0.77
AlphaEdit Fang et al. (2025) 0.98 0.89 1.00 0.96 0.93 0.85 0.98 0.92 0.91 0.79 0.94 0.88 0.84 0.77 0.56 0.72
@@
ROME Meng et al. (2022) 0.79 0.77 0.98 0.85 0.58 0.57 0.75 0.63 0.05 0.05 0.02 0.04 0.04 0.04 0.02 0.03
MEMIT Meng et al. (2023) 0.81 0.79 0.99 0.86 0.46 0.45 0.61 0.51 0.00 0.00 0.01 0.00 0.04 0.04 0.02 0.03
DEFER Hartvigsen et al. (2023) 0.59 0.68 0.79 0.69 0.44 0.43 1.00 0.62 0.40 0.40 1.00 0.60 0.40 0.39 1.00 0.60
GRACE Hartvigsen et al. (2023) 1.00 0.36 1.00 0.79 1.00 0.15 1.00 0.72 1.00 0.15 1.00 0.72 1.00 0.02 1.00 0.67
WISE Wang et al. (2024) 0.98 0.97 1.00 0.98 0.92 0.89 1.00 0.94 0.87 0.80 1.00 0.89 0.70 0.67 1.00 0.79
AlphaEdit Fang et al. (2025) 0.83 0.77 1.00 0.87 0.87 0.75 0.99 0.87 0.86 0.74 0.95 0.85 0.85 0.72 0.68 0.75
@@
ROME Meng et al. (2022) 1.52 0.97 2.02e1 0.61 1.26e5 0.02 4.07e4 0.02     2.04 0.99 3.45 0.92 1.04e2 0.03 2.41e2 0.01
MEMIT Meng et al. (2023) 1.30 0.99 4.05e2 0.93 4.56e10 0.01 2.95e6 0.00     1.64 1.00 1.59e1 0.89 9.72e1 0.04 1.32e2 0.02
DEFER Hartvigsen et al. (2023) 5.69e2 0.83 2.36e2 0.82 7.02e1 1.00 6.85e1 1.00     1.46e1 0.94 3.22e1 0.99 3.22e1 0.99 3.22e1 1.00
GRACE Hartvigsen et al. (2023) 1.05 1.00 7.10e1 1.00 7.12e1 1.00 7.73e1 1.00     1.39 1.00 5.97 1.00 9.53 1.00 9.57 1.00
WISE Wang et al. (2024) 4.93e1 0.98 1.46 0.95 2.10 0.99 3.20 0.99     1.40 1.00 2.56 0.94 1.31 0.99 5.21 0.93
AlphaEdit Fang et al. (2025) 1.58 1.00 3.12 0.98 5.97 0.93 8.49e3 0.05     1.75 1.00 1.76 1.00 2.87 0.98 1.70e2 0.88
@@
We follow prior work and evaluate the effectiveness of different editing methods based on three metrics (Wang et al., 2024): reliability, generalization and localization, which are respectively computed by model’s performance on edit, rephrased edit and irrelevant samples.
@@ 4.2 Main Experimental Results
Table 1 presents the results on the ZsRE Levy et al. (2017) dataset for the question answering task, evaluated across a range of total edits from $1$ to $1000$. ZsRE is a context-free Q&A benchmark designed to assess a model’s ability to store and retrieve specific factual knowledge. We observe that MEMOIR consistently achieves the most balanced performance across all evaluation settings. On LLaMA-3, MEMOIR maintains a reliability of 0.94 and a generalization of 0.85, with its locality reaching 1.0 across different edit counts. In contrast, other parametric methods exhibit substantial degradation as the number of edits increases. For example, WISE drops from an average metric of 0.92 to 0.77 as the number of edits grows from $1$ to 1,0001{,}000. A similar trend is observed with AlphaEdit, which declines from 0.96 to 0.72. Although GRACE, a non-parametric baseline, maintains high reliability, it suffers from poor generalization, leading to a significantly lower average metric overall compared to MEMOIR. In summary, MEMOIR offers the best trade-off between reliability, generalization, and locality, achieving an average metric of 0.93 on LLaMA-3 with 1,0001{,}000 total edits—outperforming all prior methods by a substantial margin of 0.14. Similar results are observed with Mistral, where MEMOIR again delivers the highest average metric, demonstrating its robustness and effectiveness across different LLMs.
@@
LLMs are prone to hallucinations—generating fluent but factually inaccurate or misleading content—which poses a significant challenge for their deployment in domains that demand high factual precision. We validate here the effectiveness of MEMOIR on correcting hallucinations on the SelfCheckGPT Manakul et al. (2023) dataset, which consists of Wikipedia-style biographies generated by GPT-3 Brown et al. (2020). The dataset is annotated for factual hallucinations, with ground-truth replacements verified against Wikipedia. We note that we report only locality and reliability using perplexity (PPL) metric, following (Wang et al., 2024), as no suitable metric for generalization exists in this setting. As shown in Table 2, MEMOIR again achieves the best balanced performance, with its advantages becoming even more significant with a larger number of edits. Under the most challenging setting of $600$ edits, MEMOIR maintains a saturated locality while delivering a perplexity metric, 61% and 77% smaller than WISE, the second-best performing approach, respectively on LLaMA-3 and on Mistral. These results further confirm MEMOIR’s robustness in sustaining model quality during long editing sequences.
@@
MEMIT Meng et al. (2023) 0.88 0.28 0.98 0.71 0.41 0.18 0.14 0.24
GRACE Hartvigsen et al. (2023) 0.97 0.28 1.00 0.75 0.97 0.28 1.00 0.75
WISE Wang et al. (2024) 0.99 0.36 0.98 0.78 0.96 0.37 1.00 0.78
AlphaEdit Fang et al. (2025) 0.89 0.27 0.99 0.72 0.85 0.27 0.96 0.69
@@
Effective model editing should generalize from structured prompts to distributionally shifted inputs, where the shift reflects increased complexity rather than domain change (Oren et al., 2019; Hewitt et al., 2024). We evaluate the performance of MEMOIR on OOD settings using the Temporal dataset Computer (2023), which consists of a prefix paired with a Wikipedia description and a model-generated description. As shown in Table 3, MEMOIR outperforms the state-of-the-art consistently at both $T=10$ and $T=75$, indicating strong adaptation to novel knowledge and effective generalization.
@@
In appendix B.5 we assess the efficacy of MEMOIR in multi-hop reasoning, a more challenging scenario. We evaluate on the RippleEdits benchmark (Cohen et al., 2023), where MEMOIR demonstrates better multi-hop reasoning generalization than all baseline methods.
@@ 4.3 Scaling with longer sequences of edits
MEMIT (Meng et al., 2023) 0.00 0.00 0.00 0.00
GRACE (Hartvigsen et al., 2023) 1.00 0.31 1.00 0.77 0.99 0.28 1.00 0.76
WISE (Wang et al., 2024) 0.59 0.57 1.00 0.72 0.43 0.41 1.00 0.61
AlphaEdit (Fang et al., 2025) 0.02 0.02 0.00 0.01 0.02 0.02 0.00 0.01
@@ 6 Related Works
Sparsity and non-overlapping activations have been widely explored in continual learning to mitigate catastrophic forgetting by localizing updates and reducing interference. Early works French (1991, 1999) introduced this idea, later extended by methods like PackNet Mallya and Lazebnik (2018) and Supermasks-in-Superposition Wortsman et al. (2020), which allocate disjoint parameter subsets per task but require task identities at inference and assume clear task boundaries. Gradient-based methods such as GPM Saha et al. (2021) and SPARCL Wang et al. (2022) improve parameter efficiency via orthogonal updates or activation sparsity. Even standard Dropout Srivastava et al. (2014) implicitly encourages non-overlapping activations Mirzadeh et al. (2020), while heterogeneous dropout across tasks aids activation sparsity Abbasi et al. (2022). However, these approaches remain confined to continual learning, often relying on irrelevant data or explicit regularization. In contrast, we tackle lifelong model editing without task supervision or boundaries, inducing non-overlapping updates by directly sparsifying input activations to a residual memory, enabling scalable, localized knowledge editing.
@@
Prior work in model editing falls into two categories: parametric and non-parametric methods. Parametric methods typically modify model weights through meta-learning, locating-then-editing strategies, or auxiliary memory modules. Meta-learning approaches like KE (Zheng et al., 2023) and MEND (Mitchell et al., 2022a) use hypernetworks to incorporate updates while minimizing disruption to previous knowledge. Locating-then-editing methods such as ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023) identify and edit specific modules where facts are stored, while AlphaEdit (Fang et al., 2025) improves them by projecting edits onto the null space of preserved knowledge to avoid corrupting previous information. Auxiliary module methods, including SERAC (Mitchell et al., 2022b) and T-Patcher (Huang et al., 2023), extend the model with lightweight components for storing and routing updates. WISE (Wang et al., 2024) further advances this line by introducing routing mechanisms and memory sharding to localize edits and mitigate cross-edit interference. Despite their effectiveness, parametric methods often suffer from forgetting over long edit sequences. Non-parametric methods store knowledge externally to preserve original weights. GRACE (Hartvigsen et al., 2023) retrieves discrete codebook entries to overwrite intermediate activations during inference to enable precise and local edits, while LOKA (Zhang et al., 2025) tackles simultaneous editing and unlearning by storing updates in task-specific or shared memory. However, these methods rely on exact input matches, limiting generalization to paraphrased queries.
@@ 7 Conclusion and Limitations
We introduced MEMOIR, a scalable framework for lifelong model editing that balances reliability, generalization, and locality. By sparsifying input activations through sample-dependent masks and confining updates to a dedicated parameter space, MEMOIR enables precise knowledge injection without disrupting the pre-trained model’s capabilities. At inference, it retrieves relevant updates by comparing sparse activation patterns, allowing edits to generalize to rephrased queries while preserving the model’s behavior on unrelated prompts. Extensive experiments show that MEMOIR outperforms state-of-the-art methods across question answering, hallucination correction, and out-of-distribution generalization tasks, scaling to thousands of edits with minimal forgetting. Overall, MEMOIR offers a practical tool for improving the factual accuracy and safety of deployed language models.